# RingMo-Aerial: An Aerial Remote Sensing Foundation Model With A Affine Transformation Contrastive Learning

### v1

@@ Abstract
Aerial Remote Sensing (ARS) vision tasks pose significant challenges due to the unique characteristics of their viewing angles. Existing research has primarily focused on algorithms for specific tasks, which have limited applicability in a broad range of ARS vision applications. This paper proposes the RingMo-Aerial model, aiming to fill the gap in foundation model research in the field of ARS vision. By introducing the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) mechanism and an affine transformation-based contrastive learning pre-training method, the model’s detection capability for small targets is enhanced and optimized for the tilted viewing angles characteristic of ARS. Furthermore, the ARS-Adapter, an efficient parameter fine-tuning method, is proposed to improve the model’s adaptability and effectiveness in various ARS vision tasks. Experimental results demonstrate that RingMo-Aerial achieves SOTA performance on multiple downstream tasks. This indicates the practicality and effectiveness of RingMo-Aerial in enhancing the performance of ARS vision tasks.
@@
Aerial Remote Sensing (ARS), foundation model, contrastive learning(CL), pretraining, self-supervised.
I Introduction
@@
Figure 1: Number of Aerospace Remote Sensing and Aerial Remote Sensing papers in Web of Science in the last decade. The green bars represent aerospace remote sensing, while the orange bars represent aerial remote sensing. Although research on aerial remote sensing has grown, the number of papers remains significantly lower than that of aerospace remote sensing.
@@
Figure 2: Illustration of multi-view, multi-resolution and occlusion of aerial remote sensing images:: (a) Multi-view imagery shows images captured from different angles, including tilt, top, and flat views; (b) Multi-resolution imagery demonstrates how varying low, medium, and high resolutions affect the level of detail; (c) Target occlusion compares cases where the target is partially occluded versus not occluded. These characteristics highlight the complexity of aerial remote sensing imagery in different scenarios.
Figure 3: Schematic Comparison of Aerospace Remote Sensing and Aerial Remote Sensing Image Features. Satellite images contain more low-frequency information with continuous features, while aerial images have more detailed variations and relatively less low-frequency information.
@@
Aerial Remote Sensing (ARS) is a significant method of information acquisition that, unlike satellite remote sensing, is typically not constrained by revisit cycles, satellite sensor incidence angles, or other observation timing and viewing angle limitations. Aerial remote sensing sensors are commonly mounted on Unmanned Aerial Vehicles (UAVs), airplanes, or balloons. Compared to remote sensing satellites, these aerial platforms can swiftly reach the observation area and conduct comprehensive observations from multiple perspectives and at various resolutions, thereby capturing real-time and detailed information about the target area. Consequently, aerial remote sensing finds extensive applications in emergency response, environmental monitoring, and national defense security. Therefore, research on enhancing aerial remote-sensing images’ intelligent information extraction capabilities holds significant academic and practical value.
@@
However, despite the advantages and value of aerial remote sensing imagery, our research reveals that the current focus in intelligent remote sensing interpretation remains predominantly on space-borne remote sensing. As shown in Figure 1, although research on aerial remote sensing has seen rapid growth in recent years, studies related to its intelligent interpretation remain relatively scarce. The number of published papers in this area accounts for less than 10% of space-borne remote sensing. Compared to space-borne remote sensing, aerial remote sensing possesses distinct physical characteristics, such as multi-view, multi-resolution, and occlusion effects, owing to its unique observational perspectives and sensor properties, as illustrated in Figure 2. In Figure 2(a), various perspectives are depicted, including oblique views (i and ii), a vertical view (iii), and a horizontal view (iv). Unlike the predominantly vertical perspective of satellite remote sensing, aerial remote sensing imagery captures surface information from oblique angles, enhancing the three-dimensional perception of the targets. However, this also results in distant objects appearing smaller and more densely arranged, increasing the complexity of image interpretation. Figure 2(b) compares images of different resolutions: low-resolution (i, iv), medium-resolution (ii, v), and high-resolution (iii, vi). These multi-resolution characteristics result from the varying data resolutions of aerial sensors operating at different flight altitudes and speeds. The variation in resolution affects the clarity and detail representation of the same target across different images, thereby placing higher demands on the adaptability of image interpretation models. Furthermore, aerial imagery is often characterized by significant occlusion, as shown in Figure 2(c). In this Figure, some targets are partially or fully occluded by other objects (i and iii illustrate occluded scenarios, while ii and iv show non-occluded scenarios). Such occlusions further complicate the interpretation of ARS images.
@@
Given the challenges mentioned above, it is difficult to design specialized intelligent models for aerial remote sensing that are tailored to different applications. Existing foundational models in computer vision and remote sensing vision are no longer sufficient to meet the demands of aerial remote sensing applications. The recently developed foundational models, which have gained significant attention in remote sensing interpretation, are considered an effective solution to these challenges. These models benefit from pre-training on vast amounts of data and large-scale parameters, enabling them to adapt more rapidly and effectively to various changes.
@@
However, existing foundational model studies in the remote sensing domain primarily focus on the interpretation of spaceborne remote sensing imagery, as evidenced by studies like RingMo[1] and Skysense[2]. This underscores the pressing need to develop a pre-trained foundational model specifically tailored for aerial remote sensing to address its unique application requirements. When adapting current foundational models to aerial remote sensing, several critical challenges persist:
@@
(1) Issues in the Pre-training Phase: The pre-training phase typically relies on large-scale datasets for self-supervised learning, primarily adopting two paradigms: Masked Image Modeling (MIM) and Contrastive Learning (CL). MIM, as exemplified by models like MAE [3] and SimMIM [4], involves applying masks to images and training the model to reconstruct the masked parts, thereby enhancing the model’s feature extraction capabilities. On the other hand, CL, as seen in models like MOCO [5], improves feature extraction by creating positive and negative sample pairs and training the model to learn the relationships between these pairs. Both paradigms have advantages and limitations, with MIM being more widely utilized in most visual scenarios. For instance, Swin Transformer and RingMo both utilize MIM-based pre-training methods. However, for aerial remote sensing imagery, the characteristics of images at different perspectives and resolutions differ significantly at the feature level. As shown in Figure 3, vertical images tend to exhibit more structured and continuous features, often characterized by large areas of color and texture consistency, with fewer variations in detail. Such images typically contain abundant low-frequency information due to their broad display range and relatively uniform visual presentation. In contrast, oblique images capture the sides of buildings and their height information, offering a more dynamic and three-dimensional perspective. These images are rich in detail variation and contain relatively less low-frequency information. The current self-supervised learning methods rely on reconstructing local information in images and struggle to capture these variations effectively.
(2) Issues in the Fine-tuning Phase: The traditional approach of fine-tuning the entire model presents significant challenges, primarily due to the time-consuming nature of the process. Additionally, since the foundational model is pre-trained on a diverse set of downstream tasks, fine-tuning the model globally for a specific downstream task can lead to overfitting, especially in aerial remote sensing. Unlike the single perspective and resolution typically encountered in spaceborne remote sensing, aerial remote sensing models are more prone to overfitting to specific perspectives and resolutions of a particular downstream task. This overfitting can reduce the ability of the model to generalize to different perspectives and resolutions in various downstream tasks.
@@
To address these challenges, this study introduces RingMo-Aerial, the world’s first foundational large model designed for aerial remote sensing tasks, based on the existing spaceborne remote sensing foundational model, RingMo. In response to the unique characteristics and difficulties of aerial remote sensing tasks, we propose several novel theories and methods, as shown in Figure4.
@@
First, we introduce Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) to address the multi-scale and occlusion issues caused by the oblique perspective in aerial remote sensing images. Unlike traditional multi-head self-attention mechanisms, FE-MSA incorporates an additional image patch expansion layer, ensuring that dense small objects are captured by more tokens when patches are divided into tokens for the Transformer model. Furthermore, FE-MSA leverages the Swin Transformer’s window attention and shifted window mechanisms. These improvements significantly enhance the model’s ability to detect dense small objects and occluded targets.
Second, to solve the problem of detecting and recognizing small distant targets in oblique perspectives, we propose a Contrastive Learning Affine Framework (CLAF) that scales and enlarges distant targets through affine transformations to construct positive samples. By inputting the transformed and original images as sample pairs into the model, this method guides the model to learn the unique visual characteristics of ARS images, improving its ability to recognize small targets at a distance.
@@
In addition, during the fine-tuning phase, we propose an efficient fine-tuning method for basic aerial remote sensing foundational models, called ARS-Adapter. Unlike traditional comprehensive fine-tuning methods, ARS-Adapter achieves efficient model fine-tuning by adjusting fewer than 5% of the parameters, significantly improving the training efficiency of downstream tasks while reducing the risk of overfitting. ARS-Adapter applies to all layers of the model: in the attention layers, it employs a traditional Adapter [6] structure that operates parallel to the attention layers; in the Multi-Layer Perceptron (MLP) layers, it utilizes a bottleneck structure that includes MLP downsampling, deep convolutional feature extraction, and MLP upsampling. This approach optimizes the performance of the aerial remote sensing foundational model, making it more adaptable and effective across various aerial remote sensing vision tasks.
In conclusion, the core contributions of this paper are as follows:
  • Addressing the Characteristics and Challenges of Aerial Remote Sensing Imagery: This paper introduces RingMo-Aerial, the world’s first foundational model designed for aerial remote sensing imagery. RingMo-Aerial supports various tasks across different perspectives, resolutions, and modalities, including scene classification, semantic segmentation, object detection, object tracking, change detection, and 3D reconstruction. It can be widely applied across multiple different scenarios.
@@
  • FEMSA: By incorporating a patch expansion layer and integrating the window attention mechanism, FEMSA effectively enhances the detection capability of small, dense targets in aerial remote sensing images.
  • CLAF: Utilizing affine transformations, CLAF guides the model to learn the features of aerial remote sensing imagery captured from oblique perspectives, thereby improving the model’s understanding of distant regions in oblique images.
  • ARS-Adapter: The ARS-Adapter innovatively introduces a fine-tuning method tailored to foundational models in aerial remote sensing, reducing the number of parameters adjusted, improving fine-tuning efficiency, and significantly lowering the risk of overfitting.
@@
In order to evaluate the effectiveness of the model, we performed six downstream tasks using 13 aerial remote sensing datasets and 5 aerospace remote sensing datasets with different viewpoints, resolutions and modes. The test results show that RingMo-Aerial achieves SOTA performance on all aerial remote sensing datasets and reaches the leading level on aerospace remote sensing datasets, which proves its strong generalization ability for intelligent interpretation of aerial and aerospace remote sensing images. Compared with the single-task model, RingMo-Aerial has obvious advantages in handling multi-tasks in different scenarios.
II Related Works
II-A Aerial remote sensing
@@
Aerial remote sensing offers unique characteristics and has found widespread applications across various domains. Common tasks in aerial remote sensing include object detection, object tracking, semantic segmentation, and 3D reconstruction. The most frequently utilized modalities are optical imaging [7], infrared imaging [8], and synthetic aperture radar (SAR) [9].
Compared to satellite remote sensing, aerial remote sensing provides higher image resolution and more flexible spatiotemporal coverage. However, it also presents distinct characteristics that set it apart from satellite-based methods. One notable difference is that aerial imagery is typically captured at oblique angles with varying shooting distances. This results in significant variations in the size of similar targets and increases the likelihood of target occlusion. [10, 11, 12, 13] have specifically modeled occluded targets at the structural level, significantly improving the detection and segmentation performance of occluded instances. [14, 15, 16, 17] have focused on optimizing the oblique perspective through multi-scale transformations, enhancing the resolution of small targets. Another key distinction lies in the frequency content of aerial imagery, which contains both high- and low-frequency signals, unlike satellite imagery that predominantly captures low-frequency information. Effectively leveraging high-frequency signals can greatly aid in the interpretation of aerial images. [18, 19] introduced frequency-domain enhancement modules to balance high- and low-frequency signals, leading to improved performance across various aerial remote sensing tasks.
II-B Visual foundation model
Figure 4: Comparison of model frameworks between RingMo and RingMo-Aerial.
@@ II-B Visual foundation model
Visual foundation models have gained attention for their ability to serve as general-purpose solutions in various computer vision tasks. Convolutional neural networks (CNNs) were among the earliest architectures used for visual foundation models, with well-known examples such as VGG [20] and ResNet [21]. These models rely on convolutional layers for feature extraction and are typically pre-trained on classification tasks using the ImageNet [22] dataset. With the advent of Transformers in natural language processing, visual Transformer models like ViT [23] and Swin Transformer [24] have gained prominence, utilizing masked image modeling (MIM) or contrastive learning (CL) for self-supervised pre-training. In the field of remote sensing (RS), models such as RingMo [1] and SkySense [2] have adopted the visual Transformer architecture and self-supervised pre-training strategies. Additionally, a number of novel architectures, such as ConvNeXt [25] and Mamba[26], have been proposed. However, due to technical immaturity, these newer models have yet to achieve widespread adoption.
@@
The key issue with these models lies in the pre-training phase, where they learn broad visual representations from large datasets. This chapter will also summarize three types of self-supervised visual training methods: CL, MIM, and their combination, with each method addressing both natural and RS.
II-B1 Contrastive Learning
@@
Contrastive learning is an self-supervised learning approach that emphasizes learning the common features among similar instances while distinguishing differences between dissimilar instances. Major contrastive learning methods include the MoCo [5, 27, 28] series. MoCov2 [27], an optimization of MoCo [5] inspired by SimCLR [29], introduces enhancements such as the MLP projection head and increased data augmentation. MoCov3 [28] integrates MoCov2 and SimSiam [30], but training with ViT [23] as the backbone often leads to instability and reduced model accuracy. Stability and performance are improved by freezing the patch projection layer of ViT during training. DINO [31], another self-supervised learning method, employs knowledge distillation techniques to train ViT for semantic segmentation and classification feature learning without labels, leveraging momentum encoders and multi-crop training for enhanced model performance.
@@
In the field of RS, contrastive learning also has many applications. GLCNet [32] presents a self-supervised contrastive learning method for semantic segmentation of high-resolution RS images. SeCo [33] introduces a self-supervised pretraining method that leverages temporal information in RS images to generate positive sample pairs, learning feature representations invariant and diverse to seasonal variations.
II-B2 Masked Image Modeling

### v3

# RingMo-Aerial: An Aerial Remote Sensing Foundation Model With Affine Transformation Contrastive Learning

@@ Abstract
Aerial Remote Sensing (ARS) vision tasks pose significant challenges due to the unique characteristics of their viewing angles. Existing research has primarily focused on algorithms for specific tasks, which have limited applicability in a broad range of ARS vision applications. This paper proposes the RingMo-Aerial model, aiming to fill the gap in foundation model research in the field of ARS vision. By introducing the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) mechanism and an affine transformation-based contrastive learning pre-training method, the model’s detection capability for small targets is enhanced and optimized for the tilted viewing angles characteristic of ARS. Furthermore, the ARS-Adapter, an efficient parameter fine-tuning method, is proposed to improve the model’s adaptability and effectiveness in various ARS vision tasks. Experimental results demonstrate that RingMo-Aerial achieves SOTA performance on multiple downstream tasks. This indicates the practicality and efficacy of RingMo-Aerial in enhancing the performance of ARS vision tasks.
@@
Aerial Remote Sensing (ARS), Foundation Model, Contrastive Learning(CL), Affine Transformation
@@
Figure 1: Number of Aerospace Remote Sensing and Aerial Remote Sensing papers in Web of Science in the last decade. The green bars represent aerospace remote sensing, while the orange bars represent aerial remote sensing. Although research on ARS has grown, the number of papers remains significantly lower than that of aerospace remote sensing.
@@
Figure 2: Illustration of multi-view, multi-resolution, and occlusion of ARS images: (a) Multi-view imagery shows images captured from different angles, including tilt, top, and flat views; (b) Multi-resolution imagery demonstrates how varying low, medium, and high resolutions affect the level of detail; (c) Target occlusion compares cases where the target is partially occluded versus not occluded. These characteristics highlight the complexity of ARS imagery in different scenarios.
1 Introduction
@@
Aerial Remote Sensing (ARS) is an essential method of data acquisition, distinct from aerospace remote sensing (ApRS) in its flexibility, as it is typically free from constraints such as revisit cycles, satellite sensor incidence angles, and timing limitations. ARS sensors, usually mounted on Unmanned Aerial Vehicles (UAVs), airplanes, or balloons, offer advantages over satellite-based remote sensing by quickly reaching target areas and capturing real-time, multi-perspective observations at various resolutions. This capability enables ARS to provide detailed information in applications ranging from emergency response to environmental monitoring and national defense. Thus, enhancing intelligent information extraction from ARS images holds significant academic and practical value.
Despite ARS’s distinct advantages and applications, research efforts in intelligent remote sensing interpretation are still largely focused on ApRS imagery. As shown in Fig.1, although interest in ARS has grown rapidly in recent years, studies in its intelligent interpretation remain relatively limited, with publications in this area comprising less than 10% of ApRS-related work.
@@
Unlike ApRS, ARS imagery is characterized by unique physical properties, including multi-view perspectives, varied resolutions, and occlusion effects, due to its flexible observational angles and sensor characteristics, as shown in Fig.2. In Fig.2(a), various perspectives are depicted, including oblique views (i and ii), a vertical view (iii), and a horizontal view (iv). Unlike ApRS, which predominantly captures vertical views, ARS imagery often incorporates oblique angles, providing more dimensional perspectives of targets. However, this also results in distant objects appearing smaller and more densely arranged, increasing the complexity of image interpretation. Fig.2(b) shows images at low, medium, and high resolutions (i, ii, iii, respectively), reflecting the diverse resolutions produced by ARS sensors at varying altitudes and speeds. Such differences in resolution affect target clarity and detail across images, placing higher demands on model adaptability. Additionally, significant occlusion is common in aerial imagery, as demonstrated in Fig.2(c), where some targets are partially or fully obscured by other objects, further complicating image interpretation.
@@
These challenges highlight the need for specialized intelligent models tailored to ARS applications. While recent foundational models in computer vision and RS have shown potential, they are often inadequate for ARS due to its unique requirements. These models benefit from pre-training on vast amounts of data and large-scale parameters, enabling them to adapt more rapidly and effectively to various changes.
@@
However, current foundational model research primarily focuses on ApRS imagery, as evidenced by studies like RingMo [1] and Skysense [2]. This underscores the need for a dedicated pre-trained foundational model specifically designed for ARS. Adapting existing foundational models for ARS involves addressing several critical challenges:
@@
(1) Challenges in Pre-training: Pre-training generally relies on large-scale datasets for self-supervised learning, primarily using Masked Image Modeling (MIM) and Contrastive Learning (CL). MIM, employed by models like MAE [3] and SimMIM [4], masks parts of images for reconstruction to enhance feature extraction, while CL, seen in models like MOCO [5], leverages positive and negative sample pairs to learn relational features. Both methods have merits but also limitations; MIM is widely used across visual tasks, as RingMo[1]. However, ARS images vary significantly in features due to differences in perspective and resolution. As shown in Fig.3, the vertical view displays structured, continuous features with broad areas of color consistency, containing abundant low-frequency information. In contrast, the oblique view, capturing building sides and heights, offers dynamic perspectives with high detail variability and lower low-frequency content. Current methods, which often focus on reconstructing local image information, struggle to capture these unique variations in ARS imagery.
(2) Challenges in Fine-tuning: The traditional approach of global fine-tuning of models faces significant challenges, mainly due to the fact that fine-tuning may lead to overfitting, especially in ARS. Unlike single viewpoints and resolutions that are common in spatial remote sensing, ARS models are more likely to be overfitted with specific viewpoints and resolutions for specific downstream tasks. This overfitting reduces the ability of the model to generalize to different viewpoints and resolutions in various downstream tasks.
Figure 3: Schematic Comparison of ApRS and ARS Image Features. Satellite images contain more low-frequency information with continuous features, while aerial images have more detailed variations and relatively less low-frequency information.
Figure 4: Comparison of model frameworks between RingMo and RingMo-Aerial. (a) RingMo uses SimMIM for pre-training to perform image reconstruction. (b) RingMo-Aerial builds upon this by introducing affine transformation and contrastive learning, enhancing its ability to process multi-view images. In the fine-tuning stage, the ARS adapter is introduced to adjust a small number of parameters, further enhancing the model’s generalization ability for downstream tasks.
@@
To address the challenges above, we introduces RingMo-Aerial, the first foundational large model explicitly designed for ARS tasks. Building on the existing ApRS foundational model RingMo, RingMo-Aerial addresses the unique characteristics of ARS tasks through several novel approaches, as illustrated in Fig.4.
@@
First, we propose the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) module, designed to address the multi-scale and occlusion challenges inherent to oblique perspectives in ARS imagery. Unlike traditional multi-head attention mechanisms, FE-MSA includes an image patch expansion layer to capture dense, small objects as tokens for the Transformer model, leveraging Swin Transformer’s window and shifted window attention. This enhances the model’s ability to detect densely packed and occluded objects.
Second, to improve detection and recognition of distant small targets in oblique views, we propose the Contrastive Learning Affine Framework (CLAF). CLAF uses affine transformations to scale and enlarge distant targets, constructing positive samples to guide the model in learning ARS-specific visual features.
@@
In addition, during the fine-tuning phase, we propose an efficient fine-tuning method for basic ARS foundational models called ARS-Adapter. Unlike traditional comprehensive fine-tuning methods, ARS-Adapter achieves efficient model fine-tuning by adjusting fewer than 5% of the parameters, significantly improving the training efficiency of downstream tasks while reducing the risk of overfitting. ARS-Adapter applies to all layers of the model: in the attention layers, it employs a traditional Adapter[6] structure that operates parallel to the attention layers; in the Multi-Layer Perceptron (MLP) layers, it utilizes a bottleneck structure that includes MLP downsampling, deep convolutional feature extraction, and MLP upsampling. This approach optimizes the performance of the ARS foundational model, making it more adaptable and effective across various ARS vision tasks.
In summary, the contributions of this paper can be summarised as follows:
  • Addressing ARS-specific Characteristics: We introduce RingMo-Aerial, the first foundational model for ARS, supporting multi-perspective, multi-resolution, and multi-modality tasks across various applications.
@@
  • FE-MSA Module: By incorporating a patch expansion layer and integrating the window attention mechanism, FE-MSA effectively enhances the detection capability of small, dense targets in ARS images.
  • CLAF Module: CLAF uses affine transformations to capture ARS-specific visual features, particularly in oblique perspectives, improving the model’s recognition of distant objects.
  • ARS-Adapter: The ARS-Adapter provides a fine-tuning approach tailored to ARS models, enhancing efficiency and reducing overfitting risks.
@@
To evaluate the effectiveness of the model, we performed six downstream tasks using 13 ARS datasets and 5 ApRS datasets with different viewpoints, resolutions, and modes. The test results show that RingMo-Aerial achieves SOTA performance on all ARS datasets and reaches the leading level on ApRS datasets, which proves its strong generalization ability for intelligent interpretation of aerial and ApRS images. Compared with the single-task model, RingMo-Aerial has apparent advantages in handling multi-tasks in different scenarios.
2 Related Work
2.1 Aerial remote sensing(ARS)
@@
ARS offers unique characteristics and applications across various domains, supporting tasks such as object detection, object tracking, semantic segmentation, and 3D-reconstruction. The most frequently utilized modalities are optical imaging [7], infrared imaging [8], and synthetic aperture radar (SAR) [9].
Compared to ApRS, ARS enables higher image resolution and flexible spatiotemporal coverage. However, it also presents distinct characteristics. Aerial imagery is typically captured at oblique angles with varying distances, leading to differences in target sizes and an increased likelihood of occlusion. For example, [10, 11, 12, 13] addressed target occlusion by modeling occluded objects at the structural level, significantly enhancing detection and segmentation performance. Studies such as [14, 15, 16, 17] have improved small target resolution by optimizing for oblique perspectives through multi-scale transformations.
Another critical difference lies in the frequency content of ARS imagery. Unlike ApRS images, which primarily capture low-frequency information, aerial images contain a mix of high-frequency and low-frequency signals, enabling detailed image interpretation. Recent work [18, 19] introduced frequency-domain enhancement modules to balance these signals, thereby improving performance across various ARS tasks.
2.2 Visual foundation model
@@
Visual foundation models have gained attention as general-purpose solutions for computer vision tasks. Early models, based on convolutional neural networks (CNNs) like VGG [20] and ResNet [21], used convolutional layers for feature extraction and were typically pre-trained on classification tasks using ImageNet [22]. With the advent of Transformers in natural language processing, visual Transformer models like ViT [23], and Swin Transformer [24] have gained prominence, utilizing masked image modeling (MIM) or contrastive learning (CL) for self-supervised pre-training. In the field of remote sensing (RS), models such as RingMo [1] and SkySense [2] have adopted the visual Transformer architecture and self-supervised pre-training strategies. Several novel architectures, such as ConvNeXt [25] and Mamba[26], have been proposed. However, these newer models have yet to achieve widespread adoption due to technical immaturity.
@@
A major challenge for these models lies in the pre-training phase, where they learn broad visual representations from large datasets. This section provides an overview of three self-supervised training methods—contrastive learning, masked image modeling, and their combination—each addressing both natural and RS images.
2.2.1 Contrastive Learning
@@
Contrastive learning is a self-supervised learning approach that emphasizes learning the standard features among similar instances while distinguishing differences between dissimilar cases. Primary contrastive learning methods include the MoCo [5, 27, 28] series. MoCov2 [27], an optimization of MoCo [5] inspired by SimCLR [29], introduces enhancements such as the MLP projection head and increased data augmentation. MoCov3 [28] integrates MoCov2 and SimSiam [30], but training with ViT [23] as the backbone often leads to instability and reduced model accuracy. Stability and performance are improved by freezing the patch projection layer of ViT during training. DINO [31], another self-supervised learning method, employs knowledge distillation techniques to train ViT for semantic segmentation and classification feature learning without labels, leveraging momentum encoders and multi-crop training for enhanced model performance.
@@
In RS, contrastive learning also has many applications. GLCNet [32] presents a self-supervised contrastive learning method for semantic segmentation of high-resolution RS images. SeCo [33] introduces a self-supervised pre-training method that leverages temporal information in RS images to generate positive sample pairs, learning feature representations invariant and diverse to seasonal variations.
2.2.2 Masked Image Modeling